alones (alones):

@aaronq Aaaaaaa I keep on getting -1 /.\

alones (alones):

If one endpoint of a line segment is (2, 8) and the midpoint is (-4, 7), what is the other endpoint? So, \(\frac{(2+(-4)) }{ 2 } \) \(\frac{ (8+7) }{ 2}\)

alones (alones):

yea fro the first one i get -1 gosh /.<

alones (alones):

GAWD DAMN I KNWO THE FORMULA!!! But i keep on getting -1!!!!!!

rebeccaxhawaii (rebeccaxhawaii):

chill OG. c;

OpenStudy (anonymous):

x_1 +x_2 --------- = - 4 2 x_1 = 2 x_2 you need get 2 + x_2 ------- = - 4 2 2 + x_2 = -8 x_2 = - 10 so this is the x coordinate for the second end do you can calculi the y_2 value same with this above wrote ?